# MITP-25-079

### v1

@@
Abstract: We present lattice results for $f_{K}/f_{\pi}$ in the iso-symmetric limit of pure QCD (isoQCD) with $N_{f}=2+1$ flavours, along with a determination of $|V_{us}|/|V_{ud}|$ and a study on the unitarity of the first row of the Cabibbo-Kobayashi-Maskawa (CKM) matrix after introducing strong isospin-breaking and QED effects. The results obtained are based on a combination of a Wilson unitary action and the mixed-action setup introduced in [Bussone:2025wlf, Bussone:2023kag]. The combination of the two regularisations enables a more precise control over the continuum-limit extrapolation.
@@ 1 Introduction
where $\delta_{\rm EM}^{K(\pi)}$ parameterises the QED contributions to $\Gamma\left(K(\pi)\rightarrow l\overline{\nu}_{l}[\gamma])\right)$ [Marciano:1993sh, Cirigliano:2011tm]. Furthermore, one can subsequently obtain a determination of $|V_{us}|$ employing the value of $|V_{ud}|$ as given from super-allowed nuclear $\beta$-decays [ParticleDataGroup:2024cfk], and perform a unitarity test of the first row of the CKM matrix.
In the lattice, the isospin-symmetric limit of pure QCD (isoQCD) is what is most often simulated, with two mass-degenerate light quarks and vanishing electromagnetic coupling $\alpha=0$. In order to gain access to the ratio $f_{K^{\pm}}/f_{\pi^{\pm}}$ in the full theory, we compute $f_{K}/f_{\pi}$ in isoQCD and later correct for strong isospin-breaking effects. For these effects, as well as for the QED corrections in Eq. 1.1, different approaches exist in the literature [Marciano:1993sh, Cirigliano:2011tm, DiCarlo:2019thl, Boyle:2022lsi]. Though in principle it is desirable to have a fully non-perturbative determination of these corrections, we find that the final error of our results is completely dominated by the lattice uncertainty of $f_{K}/f_{\pi}$ in isoQCD, reducing thus the impact of what one uses to correct for strong isospin-breaking and QED effects.
The computation of the ratio $f_{K}/f_{\pi}$ in isoQCD is an unambiguous prediction that can be directly compared to lattice computations by other groups, as long as the input needed to define the physical point is fixed, e.g. as in the Edinburgh Consensus [FlavourLatticeAveragingGroupFLAG:2024oxs].
For our lattice setup, we employ $N_{f}=2+1$ flavours, i.e. two mass-degenerate light and one strange flavours, based on CLS ensembles [Bruno:2014jqa, Mohler:2017wnb]. Two different lattice regularisations are employed: the first is a Wilson unitary one [Bruno:2014jqa], with Wilson fermions both in the sea and the valence sectors; the second is a mixed action one [Bussone:2025wlf, Bussone:2023kag], employing Wilson fermions in the sea and Wilson twisted mass (Wtm) [Frezzotti:2000nk, Pena:2004gb] fermions in the valence. The latter was properly tuned in order to guarantee unitarity in the continuum limit by matching the pion and kaon masses in the sea and valence sectors. This ensures that the physical quark masses in both sectors agree, recovering unitarity in the continuum limit. The combination of both lattice regularisations enables a precise control of the continuum-limit extrapolation and the enhancement of the statistical precision of the final result.
@@
In order to define the physical point at which to perform the chiral interpolation of our $f_{K}/f_{\pi}$ lattice data we employ the pion and kaon pseudoscalar masses, expressed in units of the pion decay constant. This is equivalent to setting the scale with $f_{\pi}$. These three quantities are prescribed by the Edinburgh Consensus [^1] https://indico.ph.ed.ac.uk/event/257/ and FLAG 24 [FlavourLatticeAveragingGroupFLAG:2024oxs] to be
@@
in isoQCD. Using $f_{\pi}$ to set the scale, as opposed to employing a theory scale (e.g. gradient flow scales $t_{0},w_{0}$ [Luscher:2010iy, BMW:2012hcm]), offers the benefit that its physical value is already known, as evidenced by Eq. 1.2, and thus one does not rely on the continuum-limit extrapolation of one of these scales. Conversely, $f_{\pi}$ has a strong chiral dependence. However, this effect can be properly addressed in our analysis thanks to the inclusion of physical point ensembles.
@@
The estimation of statistical uncertainties has been performed using the $\Gamma$-method [Wolff:2003sm] implementation in the ADerrors package [Ramos:2018vgu].
@@ 2 Setup
In this section we review the basic features of our setup, with an emphasis on the aspects that are most relevant to the present work. We refer the reader to [Bussone:2025wlf, Bussone:2023kag] for a fully detailed discussion of our approach.
The set of lattice ensembles under study were generated by the Coordinated Lattice Simulations (CLS) initiative [Bruno:2016plf, Mohler:2017wnb]. They use the Lüscher-Weisz gauge action [Kuberski:2023zky] and non-perturbatively $\mbox{O}(a)$-improved $N_{f}=2+1$ Wilson fermions in the sea sector. Open boundary conditions (OBC) are employed in the time direction for the gauge degrees of freedom for most ensembles, in order to avoid topology freezing as the continuum limit is approached.
@@
With the input in Eq. 1.2, the physical values $\rho_{2}^{\rm ph},\;\rho_{4}^{\rm ph}$ are unambiguously determined, thereby defining the physical point at which to chirally-interpolate the $f_{K}/f_{\pi}$ data. Note that we interpolate both in $\rho_{2}$ and $\rho_{4}$. We have one ensemble (E250) lying at the physical pion mass. In Fig. 1 we show the range of values explored for these quantities along the ensembles considered in this work.
In the valence sector, the regularisation employed for the Dirac operator is either identical to that utilised in the sea sector (Wilson unitary setup), or a Wilson twisted mass [Frezzotti:2000nk, Pena:2004gb] one (mixed-action setup). In the latter case, the valence parameters were tuned in order to match the pion and kaon masses – in units of the lattice spacing – in the sea and valence sectors [Bussone:2025wlf, Bussone:2023kag]. This ensures that physical quark masses coincide in both sectors, thereby recovering unitarity of the theory in the continuum limit, as was checked in [Bussone:2025wlf] through universality tests. Furthermore, the valence sector was tuned to maximal twist by setting the valence light PCAC standard quark mass to zero. This results in automatic $\mbox{O}(a)$-improvement up to residual sea cutoff effects of $\mbox{O}\left(ag_{0}^{4}{\rm tr}M_{q}^{\rm sea}\right)$. For a more thorough exposition about the tuning of the mixed action, see [Bussone:2025wlf, Bussone:2023kag].
@@
The relevant lattice observables for this study are the pseudoscalar masses $m_{\rm PS}$, employed to define the physical point, and decay constants $f_{\rm PS}$, with ${\rm PS}=\pi,K$. Both set of observables can be extracted from two-point functions involving the pseudoscalar density and axial current. They exhibit clear plateaus, which we extract by model averaging different plateau ranges as proposed in [Neil:2023pgt]. For more details on this latter subject, we refer to [Bussone:2025wlf, Bussone:2023kag].
@@
All physical observables are free of $\mbox{O}(a)$ cutoff effects, either thanks to the automatic $\mbox{O}(a)$-improvement feature of our mixed action, or to the inclusion of the relevant improvement coefficients for the Wilson unitary setup. Once again we refer to [Bussone:2025wlf, Bussone:2023kag] for details on improvement, as well as for the relevant renormalisation factors for the decay constants.
@@
Table 1: Set of $N_{f}=2+1$ CLS ensembles employed in this work. The table includes the values of the bare coupling $\beta=6/g_{0}^{2}$; the spatial and temporal extents of the lattice, as well as approximate values of the lattice spacing $a$ as determined in [Bussone:2025wlf]; the masses of the pion $m_{\pi}$ and the kaon $m_{K}$; and $m_{\pi}L$. Additionally, the number of configurations $N_{\rm cnfg}$ and the corresponding length of the Monte Carlo chain in Molecular Dynamics Units (MDU) employed in the computation of the mesonic correlation functions are provided. In all ensembles, open boundary conditions in the time direction are employed for the gauge fields, with the exception of ensembles E250 and D450, which have anti-periodic boundary conditions in time. The column “LMD” refers to whether the reweighting factors were computed following [Kuberski:2023zky] (“yes”) or [Luscher:2012av, Kennedy:1998cu, Clark:2003na, Clark:2006fx, Luscher:2019openQCD] (“no”). We separate by horizontal lines different ensembles and/or false replicas.
Figure 1: Measured values of $\rho_{2},\;\rho_{4}$, defined in Eq. 2.2 and Eq. 2.3, for the set of CLS ensembles employed in this work, following the ${\rm tr}(M_{q})=2m_{l}+m_{s}\approx{\rm const.}$ chiral trajectory. Empty points correspond to the Wilson unitary setup, while filled ones correspond to the Wtm mixed action. The difference in $\rho_{2}$ and $\rho_{4}$ between the Wilson and Wtm mixed action for any given ensemble is due to different cutoff effects in $f_{\pi}$ for each regularisation, as the pion and kaon masses – in lattice units – are matched to be the same in both. The physical point, marked by a black cross, is reached by interpolating both in $\rho_{2}$ and $\rho_{4}$.
@@ 3 Finite volume effects
where $m_{\eta}^{2}=\frac{4}{3}m_{K}^{2}-\frac{1}{3}m_{\pi}^{2}$, $K_{1}(x)$ is the Bessel function of the second kind, and $m(n)$ is given in [Colangelo:2005cg, Colangelo:2005gd].
The shifts in the values of $m_{\pi,K}$ and $f_{\pi,K}$ due to these contributions are well within one standard deviation, even for the smallest volume considered, corresponding to $m_{\pi}L=3.9$ in the H105 ensemble, as can be seen in Fig. 2.
@@ 4 Chiral-continuum extrapolation
We choose $\mu\equiv f_{\pi K}=\frac{2}{3}\left(f_{K}+\frac{1}{2}f_{\pi}\right)$ in the spirit of keeping it a constant renormalisation scale, since $f_{\pi K}$ has a weak chiral dependence along our ${\rm tr}\left(M_{q}\right)\approx{\rm const.}$ trajectory [Bruno:2016plf]. Furthermore, we treat the $q=f_{\pi}/f$ factor as a fit parameter, together with the LEC $L_{5}$. This amounts to ignoring the chiral dependence of $q$ due to that of $f_{\pi}$, since it enters as a higher order effect. We observe no discrepancy with this assumption when performing the chiral fits.
@@
with $\Gamma_{i}$ the various anomalous dimensions provided for our actions in [Husung:2022kvi], together with the case $\Gamma_{i}=0$. This choice for the form of cutoff effects ensures that $f_{K}/f_{\pi}$ is exactly $1$ at the SU(3) symmetric limit, even at finite lattice spacing. As probe for cutoff effects, in the spirit of not relying on any theory scale for the analysis, we will employ $af_{\pi}^{\rm sym}(\beta)$ (cf. Table LABEL:tab:fpi_sym), the pion decay constant (in lattice units) at the SU(3) symmetric point for each value of the inverse coupling $\beta$. In practice, we substitute
@@
In order to explore systematic effects in the chiral-continuum extrapolation, it is necessary to test different fit functions and cuts in the data, which we eventually average together. To this end, the first possible variation involves exploring the different available values of $\Gamma_{i}$. However, we observe no sensitivity in the data to this variation, [^2] The central value of $f_{K}/f_{\pi}$ in the continuum changes only a $0.4\%$ of $\sigma$ when varying from $\Gamma_{\rm min}=-0.11$ to $\Gamma_{\rm max}=0.76$, corresponding to the minimum possible and maximum known values of $\Gamma_{i}$ for both our Wilson unitary and Wtm mixed-action setups [Husung:2022kvi]. and thus choose to perform the final analysis only with $\Gamma_{i}=0$ in order not to explore redundant models. We note that the smallest possible value of $\Gamma_{i}$ for both our actions is $\Gamma_{i}=-0.11$.
@@
We find that expanding both A,BA,\;B in an analogous way to Eq. 4.10 always results in the fit parameter controlling the cutoff effects in $B$ being compatible with zero and thus in inflated errors, [^3] Furthermore, when employing an information criterion-based model average, e.g. employing the Takeuchi Information Criterion [Frison:2023lwb] or the Akaike Information Criterion [Neil:2023pgt], this model always results in zero weight because of this behaviour. so we only consider
@@
In order to quantify the quality of fits we employ the p-value [Bruno:2022mfy]. After selecting all models with p-value >0.1>0.1, we employ the following scheme to compute a model average result: [^5] No significant change in the final result is seen by performing a model average based on the Takeuchi Information Criterion [Frison:2023lwb] or the Akainke Information Criterion [Neil:2023pgt]. (i) compute a model average central value and statistical uncertainty with a weighted average of all models, weighting with the errors; (ii) compute a total model average uncertainty (systematic and statistical) by taking half the maximum spread between models; and (iii) considering that systematic and statistical uncertainties add up in quadrature to give the total model average uncertainty, extract the corresponding systematic uncertainty from (i) and (ii). The result is
@@
Figure 3: Model exploration for the chiral-continuum extrapolation of the ratio $f_{K}/f_{\pi}$. In the first bracket, labels $\chi\rm PT]$ and $\chi\rm PT\;ratio]$ refer to the use of Eq. 4.12 and Eq. 4.5 respectively, with cutoff effects added according to Eq. 4.11, while $f_{\pi K}$ or $f_{\pi}$ refers to what was used for the scale $\mu$ inside the chiral logarithms in Eqs. 4.7, 4.8 and 4.9. [Tay constr] refers to Eq. 4.14, and [Tay unconstr] to Eq. 4.15. The second label indicates which data are included: [-] means all data are used, $\beta>3.40]$ that the coarsest lattice spacing is removed, $m_{\pi}<360{\rm\;MeV}]$ that only ensembles with pion masses smaller than 360360 MeV are included in the analysis, and $m_{\pi}L\geq 4.2]$ the same for lattices with the volume satisfying this condition. The horizontal band in the top panel represents the final quoted result in Eq. 4.16, with statistical and systematic uncertainties added in quadrature.
Figure 4: Contribution of the different lattice ensembles to the statistical error squared in Eq. 4.16.
Figure 5: Left column: chiral-continuum extrapolation employing Eq. 4.12, using $\mu=f_{\pi}$ in the definition of the chiral logarithms, and adding cutoff effects according to Eq. 4.11 with $\Gamma_{i}=0$, removing $m_{\pi}>360$ MeV. Right column: chiral-continuum extrapolation employing Eq. 4.5, using $\mu=f_{\pi}$, and adding cutoff effects according to Eq. 4.11 with $\Gamma_{i}=0$, removing the coarsest lattice spacing $\beta=3.40$. In the two top plots, the points in the $y$-axis are projected to $\rho_{4}^{\rm ph}$ and vanishing lattice spacing, while in the two middle plots they are projected to $\rho_{2}^{\rm ph}$ and vanishing lattice spacing. The vertical dashed line marks the position of the physical point, and the bands correspond to the mass-dependence in the continuum. Finally, in the two bottom plots the data in the $y$-axis are projected to both $\rho_{2}^{\rm ph},\;\rho_{4}^{\rm ph}$. The two bands in each plot correspond to the cutoff dependence for both the Wilson unitary and Wtm mixed-actions setups. Empty points are computed with the Wilson unitary action, while filled ones are with the mixed action. The two models shown correspond to the ones giving the lowest and highest results that enter in the model average (cf. Fig. 3), and thus define the width of our error band in Eq. 4.16, corresponding to the black star point in the two bottom plots.

### v2

@@
Abstract: We present lattice results for $f_{K}/f_{\pi}$ in the iso-symmetric limit of pure QCD (isoQCD) with $N_{f}=2+1$ flavours, along with a determination of $|V_{us}|/|V_{ud}|$ and a study on the unitarity of the first row of the Cabibbo-Kobayashi-Maskawa (CKM) matrix after introducing strong isospin-breaking and QED effects. The results obtained are based on a combination of a Wilson unitary action and the mixed-action setup introduced in [12, 13]. The combination of the two regularisations enables a more precise control over the continuum-limit extrapolation.
@@ 1 Introduction
where $\delta_{\rm EM}^{K(\pi)}$ parameterises the QED contributions to $\Gamma\left(K(\pi)\rightarrow l\overline{\nu}_{l}[\gamma])\right)$ [34, 14]. Furthermore, one can subsequently obtain a determination of $|V_{us}|$ employing the value of $|V_{ud}|$ as given from super-allowed nuclear $\beta$-decays [37], and perform a unitarity test of the first row of the CKM matrix.
In the lattice, the isospin-symmetric limit of pure QCD (isoQCD) is what is most often simulated, with two mass-degenerate light quarks and vanishing electromagnetic coupling $\alpha=0$. In order to gain access to the ratio $f_{K^{\pm}}/f_{\pi^{\pm}}$ in the full theory, we compute $f_{K}/f_{\pi}$ in isoQCD and later correct for strong isospin-breaking effects. For these effects, as well as for the QED corrections in Eq. 1.1, different approaches exist in the literature [34, 14, 21, 8]. Though in principle it is desirable to have a fully non-perturbative determination of these corrections, we find that the final error of our results is completely dominated by the lattice uncertainty of $f_{K}/f_{\pi}$ in isoQCD, reducing thus the impact of what one uses to correct for strong isospin-breaking and QED effects.
The computation of the ratio $f_{K}/f_{\pi}$ in isoQCD is an unambiguous prediction that can be directly compared to lattice computations by other groups, as long as the input needed to define the physical point is fixed, e.g. as in the Edinburgh Consensus [3].
For our lattice setup, we employ $N_{f}=2+1$ flavours, i.e. two mass-degenerate light and one strange flavours, based on CLS ensembles [10, 35]. Two different lattice regularisations are employed: the first is a Wilson unitary one [10], with Wilson fermions both in the sea and the valence sectors; the second is a mixed action one [12, 13], employing Wilson fermions in the sea and Wilson twisted mass (Wtm) [25, 39] fermions in the valence. The latter was properly tuned in order to guarantee unitarity in the continuum limit by matching the pion and kaon masses in the sea and valence sectors. This ensures that the physical quark masses in both sectors agree, recovering unitarity in the continuum limit. The combination of both lattice regularisations enables a precise control of the continuum-limit extrapolation and the enhancement of the statistical precision of the final result.
@@
In order to define the physical point at which to perform the chiral interpolation of our $f_{K}/f_{\pi}$ lattice data we employ the pion and kaon pseudoscalar masses, expressed in units of the pion decay constant. This is equivalent to setting the scale with $f_{\pi}$. These three quantities are prescribed by the Edinburgh Consensus [^1] https://indico.ph.ed.ac.uk/event/257/ and FLAG 24 [3] to be
@@
in isoQCD. Using $f_{\pi}$ to set the scale, as opposed to employing a theory scale (e.g. gradient flow scales $t_{0},w_{0}$ [33, 7]), offers the benefit that its physical value is already known, as evidenced by Eq. 1.2, and thus one does not rely on the continuum-limit extrapolation of one of these scales. Conversely, $f_{\pi}$ has a strong chiral dependence. However, this effect can be properly addressed in our analysis thanks to the inclusion of physical point ensembles.
@@
The estimation of statistical uncertainties has been performed using the $\Gamma$-method [41] implementation in the ADerrors package [40].
@@ 2 Setup
In this section we review the basic features of our setup, with an emphasis on the aspects that are most relevant to the present work. We refer the reader to [12, 13] for a fully detailed discussion of our approach.
The set of lattice ensembles under study were generated by the Coordinated Lattice Simulations (CLS) initiative [9, 35]. They use the Lüscher-Weisz gauge action [30] and non-perturbatively $\mbox{O}(a)$-improved $N_{f}=2+1$ Wilson fermions in the sea sector. Open boundary conditions (OBC) are employed in the time direction for the gauge degrees of freedom for most ensembles, in order to avoid topology freezing as the continuum limit is approached.
@@
With the input in Eq. 1.2, the physical values $\rho_{2}^{\rm ph},\;\rho_{4}^{\rm ph}$ are unambiguously determined as
$\rho_{2}^{\rm ph}=1.070,\quad\rho_{4}^{\rm ph}=14.899,$ (2.4)
thereby defining the physical point at which to chirally-interpolate the $f_{K}/f_{\pi}$ data. Note that we interpolate both in $\rho_{2}$ and $\rho_{4}$. We have one ensemble (E250) lying at the physical pion mass. In Fig. 1 we show the range of values explored for these quantities along the ensembles considered in this work.
Figure 1: Measured values of $\rho_{2},\;\rho_{4}$, defined in Eq. 2.2 and Eq. 2.3, for the set of CLS ensembles employed in this work (cf. Table 1), following the ${\rm tr}(M_{q})=2m_{l}+m_{s}\approx{\rm const.}$ chiral trajectory. Empty points correspond to the Wilson unitary setup, while filled ones correspond to the Wtm mixed action. The difference in $\rho_{2}$ and $\rho_{4}$ between the Wilson and Wtm mixed action for any given ensemble is due to different cutoff effects in $f_{\pi}$ for each regularisation, as the pion and kaon masses – in lattice units – are matched to be the same in both. The physical point, marked by a black cross, is reached by interpolating both in $\rho_{2}$ and $\rho_{4}$.
In the valence sector, the regularisation employed for the Dirac operator is either identical to that utilised in the sea sector (Wilson unitary setup), or a Wilson twisted mass [25, 39] one (mixed-action setup). In the latter case, the valence parameters were tuned in order to match the pion and kaon masses – in units of the lattice spacing – in the sea and valence sectors [12, 13]. This ensures that physical quark masses coincide in both sectors, thereby recovering unitarity of the theory in the continuum limit, as was checked in [12] through universality tests. Furthermore, the valence sector was tuned to maximal twist by setting the valence light PCAC standard quark mass to zero. This results in automatic $\mbox{O}(a)$-improvement up to residual sea cutoff effects of $\mbox{O}\left(ag_{0}^{4}{\rm tr}M_{q}^{\rm sea}\right)$. For a more thorough exposition about the tuning of the mixed action, see [12, 13].
@@
The relevant lattice observables for this study are the pseudoscalar masses $m_{\rm PS}$, employed to define the physical point, and decay constants $f_{\rm PS}$, with ${\rm PS}=\pi,K$. Both set of observables can be extracted from two-point functions involving the pseudoscalar density and axial current. They exhibit clear plateaus, which we extract by model averaging different plateau ranges as proposed in [38]. For more details on this latter subject, we refer to [12, 13].
@@
All physical observables are free of $\mbox{O}(a)$ cutoff effects, either thanks to the automatic $\mbox{O}(a)$-improvement feature of our mixed action, or to the inclusion of the relevant improvement coefficients for the Wilson unitary setup. Once again we refer to [12, 13] for details on improvement, as well as for the relevant renormalisation factors for the decay constants.
@@
Table 1: Set of $N_{f}=2+1$ CLS ensembles employed in this work. We quote the values of the bare coupling $\beta=6/g_{0}^{2}$; approximate values of the lattice spacing $a$ as determined in [12]; the spatial and temporal lattice extents; the masses of the pion $m_{\pi}$ and the kaon $m_{K}$; and $m_{\pi}L$. Additionally, the number of configurations $N_{\rm cnfg}$ and the corresponding length of the Monte Carlo chain in Molecular Dynamics Units (MDU) employed in the computation of the mesonic correlation functions are provided. Open boundary conditions in the time direction are employed for the gauge fields, with the exception of ensembles E250 and D450, which have anti-periodic boundary conditions in time. The column “LMD” refers to whether the reweighting factors were computed following [30] (“yes”) or [32, 29, 16, 17, 31] (“no”). The ensemble id consists of the id used for the ensemble plus a sufix (r000,…) to distinguish different replica runs (independent Monte-Carlo runs which share the same physical and algorithmic parameters, but use different random numbers [41]). We separate by horizontal lines different ensembles. In the case of ensemble H105, we have the two replica H105r001 and H105r002. H105r005 cannot be considered a replica of H105 since different algorithmic parameters were used, and therefore it is treated as a different ensemble with the same physical parameters. The same applies to H102r001 and H102r002.
Table 2: Measurements of $\rho_{2},\;\rho_{4}$ and $f_{K}/f_{\pi}$ for both the Wilson unitary [W] and Wtm mixed-action [Wtm] setups. The mixed action was tuned in order to impose identical pion and kaon masses in the sea and valence sectors in units of the lattice spacing. The difference between $\rho_{2},\;\rho_{4}$ in the Wilson unitary and mixed-action regularisations therefore comes exclusively from different cutoff effects in $f_{\pi}$ for each regularisation. Finally, the ratio $f_{K}/f_{\pi}$ is $1.0$ without error in the SU(3) symmetric point by construction. All measurements are corrected for finite-volume effects as detailed in Section 3. Measurements corresponding to ensembles H105 and H105r005 were averaged together, as well as those of H102r001 and H102r002 (cf. Table 1). All quantities are renormalised and non-perturbatively $\mbox{O}(a)$-improved, for a general discussion we refer to [12, 13].
ID $\rho_{2}$ [W] $\rho_{2}$ [Wtm] $\rho_{4}$ [W] $\rho_{4}$ [Wtm] $f_{K}/f_{\pi}$ [W] $f_{K}/f_{\pi}$ [Wtm]
H101 $8.316(76)$ $7.235(52)$ $12.47(11)$ $10.85(8)$ $1.0$ $1.0$
H102 $6.542(56)$ $5.716(45)$ $13.35(11)$ $11.66(8)$ $1.0509(21)$ $1.0569(16)$
H105 $4.617(83)$ $3.873(49)$ $14.92(23)$ $12.51(10)$ $1.1182(78)$ $1.1021(33)$
H400 $8.474(89)$ $7.524(77)$ $12.71(13)$ $11.29(12)$ $1.0$ $1.0$
D450 $2.859(56)$ $2.503(32)$ $15.10(22)$ $13.22(18)$ $1.1418(66)$ $1.1422(33)$
N202 $7.834(60)$ $7.238(68)$ $11.75(9)$ $10.86(10)$ $1.0$ $1.0$
N203 $5.884(38)$ $5.357(47)$ $12.61(7)$ $11.48(10)$ $1.0540(15)$ $1.0519(17)$
N200 $4.351(40)$ $4.057(36)$ $13.76(9)$ $12.83(10)$ $1.0998(27)$ $1.1070(34)$
D200 $2.392(26)$ $2.142(27)$ $14.94(11)$ $13.37(12)$ $1.1551(33)$ $1.1343(46)$
E250 $1.100(43)$ $0.997(18)$ $16.41(56)$ $14.88(26)$ $1.1952(215)$ $1.1937(83)$
N300 $7.965(53)$ $7.635(60)$ $11.95(8)$ $11.45(9)$ $1.0$ $1.0$
N302 $5.668(84)$ $5.427(72)$ $12.53(16)$ $12.00(15)$ $1.0574(20)$ $1.0608(42)$
J303 $3.545(35)$ $3.326(46)$ $13.97(12)$ $13.11(16)$ $1.1188(44)$ $1.1161(64)$
E300 $1.889(38)$ $1.775(29)$ $15.79(27)$ $14.83(18)$ $1.1735(138)$ $1.1893(201)$
J500 $7.443(130)$ $7.239(165)$ $11.16(20)$ $10.86(25)$ $1.0$ $1.0$
J501 $5.374(88)$ $5.292(118)$ $12.30(19)$ $12.12(26)$ $1.0607(28)$ $1.0647(51)$
Table 3: Values for $af_{\pi}$ at the SU(3) symmetric point, defined by $m_{\pi}=m_{K}\approx 420$ MeV, for each $\beta$ considered in this work, to be employed as probes for the cutoff effects (cf. Section 4). We quote values for both the Wilson unitary [W] and Wtm mixed-action [Wtm] setups, the difference between them being a cutoff effect. Finite-volume effects are included as detailed in Section 3. Results are renormalised and non-perturbatively $\mbox{O}(a)$-improved, for a general discussion we refer to [12, 13]
$\beta$ $af_{\pi}^{\rm sym}$ [W] $af_{\pi}^{\rm sym}$ [Wtm]
$3.40$ $0.06319(28)$ $0.06774(24)$
$3.46$ $0.05620(25)$ $0.05965(30)$
$3.55$ $0.04792(17)$ $0.04985(19)$
$3.70$ $0.03770(11)$ $0.03850(13)$
$3.85$ $0.02991(23)$ $0.03033(33)$
@@ 3 Finite volume effects
where $m_{\eta}^{2}=\frac{4}{3}m_{K}^{2}-\frac{1}{3}m_{\pi}^{2}$, $K_{1}(x)$ is the Bessel function of the second kind, and $m(n)$ is given in [19, 18].
The shifts in the values of $m_{\pi,K}$ and $f_{\pi,K}$ due to the finite-volume corrections are well within our present statistical errors, even for the smallest volume considered, corresponding to $m_{\pi}L=3.9$ in the H105 ensemble, as can be seen in Fig. 2.
@@ 4 Chiral-continuum extrapolation
We choose $\mu\equiv f_{\pi K}=\frac{2}{3}\left(f_{K}+\frac{1}{2}f_{\pi}\right)$ in the spirit of keeping it a constant renormalisation scale, since $f_{\pi K}$ has a weak chiral dependence along our ${\rm tr}\left(M_{q}\right)\approx{\rm const.}$ trajectory [9]. Furthermore, we treat the $q=f_{\pi}/f$ factor as a fit parameter, together with the LEC $L_{5}$. This amounts to ignoring the chiral dependence of $q$ due to that of $f_{\pi}$, since it enters as a higher order effect. We observe no discrepancy with this assumption when performing the chiral fits.
@@
with $\Gamma_{i}$ the various anomalous dimensions provided for our actions in [28], together with the case $\Gamma_{i}=0$. This choice for the form of cutoff effects ensures that $f_{K}/f_{\pi}$ is exactly 11 at the SU(3) symmetric limit, even at finite lattice spacing. As probe for cutoff effects, in the spirit of not relying on any theory scale for the analysis, we will employ $af_{\pi}^{\rm sym}(\beta)$ (cf. Table LABEL:tab:fpi_sym), the pion decay constant (in lattice units) at the SU(3) symmetric point for each value of the inverse coupling $\beta$. In practice, we substitute
@@
In order to explore systematic effects in the chiral-continuum extrapolation, it is necessary to test different fit functions and cuts in the data, which we eventually average together. To this end, the first possible variation involves exploring the different available values of $\Gamma_{i}$. However, we observe no sensitivity in the data to this variation, [^2] The central value of $f_{K}/f_{\pi}$ in the continuum changes only a $0.4\%$ of $\sigma$ when varying from $\Gamma_{\rm min}=-0.11$ to $\Gamma_{\rm max}=0.76$, corresponding to the minimum possible and maximum known values of $\Gamma_{i}$ for both our Wilson unitary and Wtm mixed-action setups [28]. and thus choose to perform the final analysis only with $\Gamma_{i}=0$ in order not to explore redundant models. We note that the smallest possible value of $\Gamma_{i}$ for both our actions is $\Gamma_{i}=-0.11$.
@@
We find that expanding both A,BA,\;B in an analogous way to Eq. 4.10 always results in the fit parameter controlling the cutoff effects in $B$ being compatible with zero and thus in inflated errors, [^3] Furthermore, when employing an information criterion-based model average, e.g. employing the Takeuchi Information Criterion [26] or the Akaike Information Criterion [38], this model always results in zero weight because of this behaviour. so we only consider
@@
In order to quantify the quality of fits we employ the p-value [11]. After selecting all models with p-value >0.1>0.1, we employ the following scheme to compute a model average result: [^5] No significant change in the final result is seen by performing a model average based on the Takeuchi Information Criterion [26] or the Akainke Information Criterion [38]. (i) compute a model average central value and statistical uncertainty with a weighted average of all models, weighting with the errors; (ii) compute a total model average uncertainty (systematic and statistical) by taking half the maximum spread between models; and (iii) considering that systematic and statistical uncertainties add up in quadrature to give the total model average uncertainty, extract the corresponding systematic uncertainty from (i) and (ii). The result is
@@
Figure 3: Model exploration for the chiral-continuum extrapolation of the ratio $f_{K}/f_{\pi}$. In the first bracket, labels $\chi\rm PT]$ and $\chi\rm PT\;ratio]$ refer to the use of Eq. 4.12 and Eq. 4.5 respectively, with cutoff effects added according to Eq. 4.11, while $f_{\pi K}$ or $f_{\pi}$ refers to what was used for the scale $\mu$ inside the chiral logarithms in Eqs. 4.7, 4.8 and 4.9. [Tay constr] refers to Eq. 4.14, and [Tay unconstr] to Eq. 4.15. The second label indicates which data are included: [-] means all data are used, $\beta>3.40]$ that the coarsest lattice spacing is removed, $m_{\pi}<360{\rm\;MeV}]$ that only ensembles with pion masses smaller than $360$ MeV are included in the analysis, and $m_{\pi}L\geq 4.2]$ the same for lattices with the volume satisfying this condition. The horizontal band in the top panel represents the final quoted result in Eq. 4.16, with statistical and systematic uncertainties added in quadrature.
Figure 4: Contribution of the different lattice ensembles to the statistical error squared in Eq. 4.16. Large contributions indicate that the corresponding ensembles are strongly constraining the chiral-continuum extrapolation of $f_{K}/f_{\pi}$, and therefore are expected to be ensembles close to the continuum limit and physical point. In particular, ensemble J501 is the only ensemble at the finest lattice spacing ($a\approx 0.038$ fm) that lies outside the SU(3) symmetric point. Ensemble J303, on the other hand, is at the second finest lattice spacing ($a\approx 0.049$ fm) and close to the physical point ($m_{\pi}\approx 261$ MeV). Although we also have an ensemble at the same lattice spacing that is closer to the physical point (E300, $m_{\pi}\approx 177$ MeV), its statistical sample is significantly smaller (cf. Table 1 and Table LABEL:tab:measurements), which reduces its overall impact on the fits.
Figure 5: Left column: chiral-continuum extrapolation employing Eq. 4.12, using $\mu=f_{\pi}$ in the definition of the chiral logarithms, and adding cutoff effects according to Eq. 4.11 with $\Gamma_{i}=0$, removing $m_{\pi}>360$ MeV. We note that the $x$-axis direction has been reversed relative to the one in the right column in order to facilitate a direct comparison of the physical point extrapolation in both fits. Right column: chiral-continuum extrapolation employing Eq. 4.5, using $\mu=f_{\pi}$, and adding cutoff effects according to Eq. 4.11 with $\Gamma_{i}=0$, removing the coarsest lattice spacing $\beta=3.40$. In the two top plots, the points in the $y$-axis are projected to $\rho_{4}^{\rm ph}$ and vanishing lattice spacing, while in the two middle plots they are projected to $\rho_{2}^{\rm ph}$ and vanishing lattice spacing. The vertical dashed line marks the position of the physical point, and the bands correspond to the mass-dependence in the continuum. Finally, in the two bottom plots the data in the $y$-axis are projected to both $\rho_{2}^{\rm ph},\;\rho_{4}^{\rm ph}$. The two bands in each plot correspond to the cutoff dependence for both the Wilson unitary and Wtm mixed-actions setups. Empty points are computed with the Wilson unitary action, while filled ones are with the mixed action. The two models shown correspond to the ones giving the lowest and highest results that enter in the model average (cf. Fig. 3), and thus define the width of our error band in Eq. 4.16, corresponding to the black star point in the two bottom plots.